OpenStudy (anonymous):

how to simplify (6x^5)^3

OpenStudy (anonymous):

Distribute the parenthesis.

OpenStudy (anonymous):

216x^15

OpenStudy (anonymous):

Use rule: \(\huge (a^x)^y=a^{xy}\) So, \(\Large (6x^5)^3=6^3x^{15}=216x^{15} \)